# WebAgent-R1: Training Web Agents via End-to-End Multi-Turn Reinforcement Learning

### v1

@@ Abstract
While reinforcement learning (RL) has demonstrated remarkable success in enhancing large language models (LLMs), it has primarily focused on single-turn tasks such as solving math problems. Training effective web agents for multi-turn interactions remains challenging due to the complexity of long-horizon decision-making across dynamic web interfaces. In this work, we present WebAgent-R1, a simple yet effective end-to-end multi-turn RL framework for training web agents. It learns directly from online interactions with web environments by asynchronously generating diverse trajectories, entirely guided by binary rewards depending on task success. Experiments on the WebArena-Lite benchmark demonstrate the effectiveness of WebAgent-R1, boosting the task success rate of Qwen-2.5-3B from 6.1% to 33.9% and Llama-3.1-8B from 8.5% to 44.8%, significantly outperforming existing state-of-the-art methods and strong proprietary models such as OpenAI o3. In-depth analyses reveal the effectiveness of the thinking-based prompting strategy and test-time scaling through increased interactions for web tasks. We further investigate different RL initialization policies by introducing two variants, namely WebAgent-R1-Zero and WebAgent-R1-CoT, which highlight the importance of the warm-up training stage (i.e., behavior cloning) and provide insights on incorporating long chain-of-thought (CoT) reasoning in web agents. [^1] Code and artifacts are available at https://github.com/weizhepei/WebAgent-R1
@@ 1 Introduction
Meanwhile, several concurrent works have explored end-to-end RL with on-policy updates for training LLM agents in multi-turn interactive scenarios, such as simulated games and coding environments (Wang et al., 2025; Cao et al., 2025). Unlike off-policy RL that trains on data generated by older versions of the agent, on-policy RL collects training data directly from the agent’s current behavior. This ensures that the learning process is better aligned with the agent’s most recent actions, often leading to more stable and effective learning (Schulman et al., 2015, 2017). It also eliminates the need for additional overheads in off-policy RL (e.g., maintaining a replay buffer and filtering outdated trajectories), and enables the agent to behave adaptively based on its own past decisions—a key advantage in interactive environments where early decisions can significantly affect next steps.
@@
In light of this, we propose WebAgent-R1, an end-to-end multi-turn RL framework for training web agents. Specifically, our design addresses several key challenges in this setting. First, at each step, the environmental observation (e.g., HTML content) can span thousands of tokens, causing the accumulated context over long horizons to incur substantial memory overheads. To mitigate this, we introduce a dynamic context compression mechanism, which adaptively adjusts the contexts across turns, ensuring scalability and preventing out-of-memory issues. Second, existing RL solutions for LLM agents are not well-suited for multi-turn scenarios. Inspired by group relative policy optimization (GRPO) (Shao et al., 2024), we extend it to multi-turn settings (M-GRPO) and employ an asynchronous trajectory rollout strategy to further improve training efficiency by generating multiple trajectories in parallel. These designs enable efficient RL training and lead to state-of-the-art performance on the WebArena-Lite benchmark, as shown in Figure 1. Extensive ablations further validate our key design choices, reveal an effective test-time scaling strategy for web tasks, and offer insights into the roles of behavior cloning and long CoT reasoning in RL-based web agent training.
@@
  • We implement an end-to-end multi-turn RL framework for training web agents, with dynamic context compression and asynchronous trajectory rollout mechanisms to achieve training efficiency.
@@ 2.1 Problem Formulation
Following prior work (Qi et al., 2025), we adopt WebArena (Zhou et al., 2024a) as the web environment over other simulated or static environments such as WebShop (Yao et al., 2022) or Mind2Web (Deng et al., 2023) for greater practicality—It provides a realistic, self-hostable environment for web agents, along with rule-based rubrics that automatically check for indicators of success in the final state (e.g., confirmation messages or expected content on the page). Note that some prior works (Liu et al., 2025; He et al., 2024a) incorporate web page screenshots as additional visual inputs, whereas our work focuses solely on text-based decision-making over HTML. Other efforts, such as Yang et al. (2025b), explore optimizing the action space or prompt design without model fine-tuning. These directions are orthogonal to our investigated problem and can be conceptually integrated with our method as future work.
@@ 2.3 End-to-End Multi-Turn Reinforcement Learning
As illustrated in Figure 2, our end-to-end multi-turn RL framework trains web agents through online interactions guided by rule-based outcome rewards. To enable efficient and scalable training, we implemented two key mechanisms: dynamic context compression to reduce memory overhead, and asynchronous trajectory rollout to improve sampling efficiency. Based on the BC-trained policy, we further fine-tune the agent using an extension of GRPO (Qi et al., 2025) in the multi-turn settings, termed M-GRPO. Our implementation can be viewed as a minimalist approach that supports efficient multi-turn RL training while maintaining generality, with potential for future extensions (e.g., incorporating fine-grained reward shaping mechanisms for intermediate steps).
@@
In web tasks, each observation $s_{t}$ often contains thousands of tokens. Across multi-turn interactions, the accumulated context grows rapidly, leading to excessive memory usage and potential out-of-memory issues, making training impractical. To address this, we propose a dynamic context compression strategy. As new observations arrive, earlier ones are simplified to reduce the context length while preserving the complete action history. Let the interaction history at step $t$ be $h_{t}=(s^{\prime}_{1},a_{1},s^{\prime}_{2},a_{2},\ldots,s_{t})$, where each $s^{\prime}_{i}$ is a simplified template (e.g., ‘‘Simplified HTML’’) representing prior observations. When the agent executes an action $a_{t}$ and receives a new observation $s_{t+1}$, the updated history becomes $h_{t+1}=(s^{\prime}_{1},a_{1},s^{\prime}_{2},a_{2},\ldots,s^{\prime}_{t},a_{t},s_{t+1})$, where $s_{t}$ is replaced by its simplified version $s^{\prime}_{t}$. This allows the agent to maintain a compact yet informative context of past interactions. Since the context evolves dynamically, we also update the loss masks accordingly to ensure that the loss is correctly computed only on the action tokens during the M-GRPO optimization.
@@
Table 2: Task success rate (SR) comparison across different methods on various websites in WebArena-Lite (Liu et al., 2025; Qi et al., 2025; Zhou et al., 2024a). Baseline performance is reported as the higher value between our reproduced results and those reported in the literature (Qi et al., 2025). The best scores are highlighted in bold.
@@
Asynchronous Trajectory Rollout
Generating a group of trajectories requires repeated interaction with the environment and can be time-consuming. To address this, we introduce an asynchronous trajectory rollout strategy, where multiple independent browser instances $\{\mathcal{E}_{1},\mathcal{E}_{2},\cdots,\mathcal{E}_{G}\}$ are instantiated, each maintaining its own context (e.g., cookies). For each task, all instances are initialized with the same starting page, but the agent interacts with them independently, resulting in diverse histories and trajectories. This asynchronous design enables efficient trajectory generation in M-GRPO.
@@
We use the default rule-based reward functions in the web environment, which assign binary rewards ($r{=}1$ for success, $r{=}0$ otherwise) based on task-specific criteria (e.g., reaching a target page). This eliminates the need for outcome reward models (Qi et al., 2025), ensuring a simple and generalizable training setup.
@@ 3.1 Experimental Setup
Like prior works (Liu et al., 2025; Qi et al., 2025), we focus on web agents for real-world scenarios, specifically utilizing WebArena (Zhou et al., 2024a), a self-hostable and realistic web environment that supports practical tasks across diverse domains: social forums (Reddit), collaborative coding (GitLab), e-commerce content management systems (CMS), open street maps (Map), and online shopping (Shopping).
@@
For prompting baselines, we provide a comprehensive comparison with both open-source and proprietary models, including general-purpose models (e.g., Qwen2.5, Llama3.1, GPT-4) and reasoning-specialized models (e.g., QwQ, OpenAI o3 (OpenAI, 2025)), covering various model sizes. For finetuning methods, we employ Qwen2.5-3B and Llama3.1-8B as the backbone model.
@@ 3.5 Analysis
As shown in Table 3, using the thinking format significantly improves task success rates across models, particularly for stronger ones (e.g., o4-mini improves from 15.9% to 36.9%). Interestingly, while the average single-turn response length remains similar (e.g., 139 $\rightarrow$ 142 tokens for Qwen2.5-3B), the number of interactions increases substantially (e.g., 6 $\rightarrow$ 17) with the thinking format. This observation suggests a novel test-time scaling strategy for web tasks—rather than producing longer single-turn responses, the web agent can become more effective by engaging in deeper multi-turn interactions.
@@
Building on the above finding, we further investigate how increasing the number of interactions between the web agent and the environment affects performance. As shown in Figure 5, allowing more interaction turns consistently improves success rates across prompting-based, SFT, and RL-based methods. We hypothesize that this form of test-time scaling facilitates deeper exploration and yields longer trajectories, potentially enabling the agent to iteratively refine its actions and make more informed decisions through extended interactions.
@@ 4.1 LLM-based Agents
LLMs have demonstrated promising agentic capabilities, such as breaking down complex tasks into manageable subgoals and reasoning over long horizons (Zhou et al., 2022; Huang et al., 2022; Madaan et al., 2022; Li et al., 2023a, b; Wu et al., 2024; Chu et al., 2025). Building on these capabilities, LLM-based agents have been applied to a variety of real-world interactive tasks, including web navigation (Nakano et al., 2021; Yao et al., 2022; Ma et al., 2023; Gur et al., 2024; Abuelsaad et al., 2024; Lutz et al., 2024; Patel et al., 2024; Putta et al., 2024), general computer use (Li et al., 2020; Deng et al., 2023; Yang et al., 2024), and embodied environments (Puig et al., 2018; Shridhar et al., 2020; Toyama et al., 2021; Fan et al., 2022; Huang et al., 2022). Specifically, our work focuses on text-based web agents that operate in browser-based environments purely based on HTML content, which requires agentic capabilities such as tool use, memory, and decision-making under partial observability (Zhou et al., 2024a; Qi et al., 2025). Complementary to this line of work, GUI agents leverage additional multimodal inputs such as screenshots, enabling visual-guided interactions with the environment (Lee et al., 2023; Shaw et al., 2023; Zheng et al., 2024; He et al., 2024a, b; Koh et al., 2024; Kil et al., 2024; Lei et al., 2025; Liu et al., 2025). For a comprehensive overview, we refer readers to recent surveys (Wang et al., 2024a; Hu et al., 2025; Ning et al., 2025).
@@ 4.2 Reinforcement Learning for LLMs
Recent advances like DeepSeek-R1 (Guo et al., 2025) highlight the strong potential of RL in enhancing LLMs. However, most prior work focuses on single-turn tasks such as math problems (Shao et al., 2024), with limited exploration in multi-turn settings (Zhou et al., 2024b, 2025). Recent efforts have made some progress in this direction, such as training LLM agents to repeatedly use search engines (Jin et al., 2025; Sun et al., 2025; Chen et al., 2025; Song et al., 2025), but typically constrain actions to simple API calls without real environment interaction. A few concurrent works, such as RAGEN (Wang et al., 2025) and SkyRL (Cao et al., 2025), have applied RL to more dynamic settings like simulated games and coding environments (Jimenez et al., 2024). However, real-world web environments remain largely underexplored. Our work fills this gap by providing a practical framework and offering actionable insights for training web agents with end-to-end RL.
@@ 5 Conclusion
This work introduces WebAgent-R1, an end-to-end multi-turn RL framework for training web agents. We extend the standard GRPO to multi-turn settings, termed M-GRPO, and implement dynamic context compression and asynchronous trajectory rollout mechanisms for efficient training. Empirically, WebAgent-R1 achieves new state-of-the-art results on the WebArena-Lite benchmark. Our findings underscore the critical role of behavior cloning in initializing web agents, providing a strong foundation for effective RL. We further analyze training dynamics and explore the effects of thinking-based prompting and test-time scaling strategies, showing that increasing interaction depth consistently enhances web agents. Future work includes exploring multi-modal inputs and extending our approach to broader GUI-based tasks beyond web environments, such as computer use.

### v2

@@ Abstract
While reinforcement learning (RL) has demonstrated remarkable success in enhancing large language models (LLMs), it has primarily focused on single-turn tasks such as solving math problems. Training effective web agents for multi-turn interactions remains challenging due to the complexity of long-horizon decision-making across dynamic web interfaces. In this work, we present WebAgent-R1, a simple yet effective end-to-end multi-turn RL framework for training web agents. It learns directly from online interactions with web environments by generating diverse trajectories in parallel, entirely guided by binary rewards depending on task success. Experiments on the WebArena-Lite benchmark demonstrate the effectiveness of WebAgent-R1, boosting the task success rate of Qwen-2.5-3B from 6.1% to 33.9% and Llama-3.1-8B from 8.5% to 44.8%, significantly outperforming existing state-of-the-art methods and strong proprietary models such as OpenAI o3. In-depth analyses reveal the effectiveness of the thinking-based prompting strategy and test-time scaling through increased interactions for web tasks. We further investigate different RL initialization policies by introducing two variants, namely WebAgent-R1-Zero and WebAgent-R1-CoT, which highlight the importance of the warm-up training stage (i.e., behavior cloning) and provide insights on incorporating long chain-of-thought (CoT) reasoning in web agents. [^1] Code and artifacts are available at https://github.com/weizhepei/WebAgent-R1
@@ 1 Introduction
Meanwhile, several concurrent works have explored end-to-end RL with on-policy updates for training LLM agents in multi-turn interactive scenarios, such as simulated games and coding environments (Wang et al., 2025b; Cao et al., 2025). Unlike off-policy RL that trains on data generated by older versions of the agent, on-policy RL collects training data directly from the agent’s current behavior. This ensures that the learning process is better aligned with the agent’s most recent actions, often leading to more stable and effective learning (Schulman et al., 2015, 2017). It also eliminates the need for additional overheads in off-policy RL (e.g., maintaining a replay buffer and filtering outdated trajectories), and enables the agent to behave adaptively based on its own past decisions—a key advantage in interactive environments where early decisions can significantly affect next steps.
@@
In light of this, we propose WebAgent-R1, an end-to-end multi-turn RL framework for training web agents. Specifically, our design addresses several key challenges in this setting. First, at each step, the environmental observation (e.g., HTML content) can span thousands of tokens, causing the accumulated context over long horizons to incur substantial memory overheads. To mitigate this, we introduce a dynamic context compression mechanism, which adaptively adjusts the contexts across turns, ensuring scalability and preventing out-of-memory issues. Second, existing RL solutions for LLM agents are not well-suited for multi-turn scenarios. Inspired by group relative policy optimization (GRPO) (Shao et al., 2024), we extend it to multi-turn settings (M-GRPO) and employ a parallel trajectory rollout strategy to further improve training efficiency by generating multiple trajectories in parallel. These designs enable efficient RL training and lead to state-of-the-art performance on the WebArena-Lite benchmark, as shown in Figure 1. Extensive ablations further validate our key design choices, reveal an effective test-time scaling strategy for web tasks, and offer insights into the roles of behavior cloning and long CoT reasoning in RL-based web agent training.
@@
  • We implement an end-to-end multi-turn RL framework for training web agents, with dynamic context compression and parallel trajectory rollout mechanisms to achieve training efficiency.
@@ 2.1 Problem Formulation
Following prior work (Qi et al., 2025), we adopt WebArena (Zhou et al., 2024a) as the web environment over other simulated or static environments such as WebShop (Yao et al., 2022) or Mind2Web (Deng et al., 2023) for greater practicality—It provides a realistic, self-hostable environment for web agents, along with rule-based rubrics that automatically check for indicators of success in the final state (e.g., confirmation messages or expected content on the page). Note that some prior works (Liu et al., 2025a; He et al., 2024a) incorporate web page screenshots as additional visual inputs, whereas our work focuses solely on text-based decision-making over HTML. Other efforts, such as Yang et al. (2025b), explore optimizing the action space or prompt design without model fine-tuning. These directions are orthogonal to our investigated problem and can be conceptually integrated with our method as future work.
@@ 2.3 End-to-End Multi-Turn Reinforcement Learning
As illustrated in Figure 2, our end-to-end multi-turn RL framework trains web agents through online interactions guided by rule-based outcome rewards. To enable efficient and scalable training, we implemented two key mechanisms: dynamic context compression to reduce memory overhead, and parallel trajectory rollout to improve sampling efficiency. Based on the BC-trained policy, we further fine-tune the agent using an extension of GRPO (Qi et al., 2025) in the multi-turn settings, termed M-GRPO. Our implementation can be viewed as a minimalist approach that supports efficient multi-turn RL training while maintaining generality, with potential for future extensions (e.g., allowing fine-grained reward shaping for intermediate steps).
@@
In web tasks, each observation $s_{t}$ often contains thousands of tokens. Across multi-turn interactions, the accumulated context grows rapidly, leading to excessive memory usage and potential out-of-memory issues, making training impractical. To address this, we propose a dynamic context compression strategy. As new observations arrive, earlier ones are simplified to reduce the context length while preserving the complete action history. Let the interaction history at step $t$ be $h_{t}=(s^{\prime}_{1},a_{1},s^{\prime}_{2},a_{2},\ldots,s_{t})$, where each $s^{\prime}_{i}$ is a template with only a few tokens (e.g., $s^{\prime}_{i}=\texttt{``Simplified HTML''}$) representing prior observations. When the agent executes an action $a_{t}$ and receives a new observation $s_{t+1}$, the updated history becomes $h_{t+1}=(s^{\prime}_{1},a_{1},s^{\prime}_{2},a_{2},\ldots,s^{\prime}_{t},a_{t},s_{t+1})$, where $s_{t}$ is replaced by its simplified version $s^{\prime}_{t}$. This allows the agent to maintain a compact yet informative context of past interactions. Since the context evolves dynamically, we also update the loss masks accordingly to ensure that the loss is correctly computed only on the action tokens during the M-GRPO optimization.
@@
Table 2: Task success rate (SR) comparison across different methods on various websites in WebArena-Lite (Liu et al., 2025a; Qi et al., 2025; Zhou et al., 2024a). Baseline performance is reported as the higher value between our reproduced results and those reported in the literature (Qi et al., 2025). The best scores are highlighted in bold.
@@
Parallel Trajectory Rollout
Generating a group of trajectories requires repeated interaction with the environment and can be time-consuming. To address this, we introduce a parallel trajectory rollout strategy, where multiple independent browser instances $\{\mathcal{E}_{1},\mathcal{E}_{2},\cdots,\mathcal{E}_{G}\}$ are instantiated, each maintaining its own context (e.g., cookies). For each task, all instances are initialized with the same starting page, but the agent interacts with them independently, resulting in diverse histories and trajectories. This parallel design enables efficient trajectory generation in M-GRPO.
@@
We use the default rule-based reward functions in the web environment, which assign binary rewards (r=1r{=}1 for success, r=0r{=}0 otherwise) based on task-specific criteria (e.g., reaching a target page). This eliminates the need for outcome reward models (Qi et al., 2025), ensuring a simple and generalizable training setup.
@@ 3.1 Experimental Setup
Like prior works (Liu et al., 2025a; Qi et al., 2025), we focus on web agents for real-world scenarios, specifically utilizing WebArena (Zhou et al., 2024a), a self-hostable and realistic web environment that supports practical tasks across diverse domains: social forums (Reddit), collaborative coding (GitLab), e-commerce content management systems (CMS), open street maps (Map), and online shopping (Shopping).
@@
For prompting baselines, we provide a comprehensive comparison with both open-source and proprietary models, including general-purpose models (e.g., Qwen2.5, Llama3.1, GPT-4) and reasoning-specialized models (e.g., QwQ, OpenAI o3 (OpenAI, 2025)), covering various model sizes. For finetuning methods, we employ Qwen2.5-3B and Llama3.1-8B as the backbone model. We refer the readers to (Liu et al., 2025a) for more baseline results on the WebArena-Lite benchmark.
@@ 3.5 Analysis
As shown in Table 3, using the thinking format significantly improves task success rates across models, particularly for stronger ones (e.g., o4-mini improves from 15.9% to 36.9%). Interestingly, while the average single-turn response length remains similar (e.g., 139 $\rightarrow$ 142 tokens for Qwen2.5-3B), the number of interactions increases substantially (e.g., 6 $\rightarrow$ 17 for Qwen2.5-3B) with the thinking format. These results indicate that prompting with explicit thinking instructions enhances web agents by encouraging more frequent interactions. As a result, we conjecture that this observation suggests a novel test-time scaling strategy for web tasks—rather than producing longer single-turn responses, the web agent can become more effective by engaging in deeper multi-turn interactions.
@@
Building on the above finding, we further investigate how increasing the number of interactions between the web agent and the environment affects performance. As shown in Figure 5, allowing more interaction turns consistently improves success rates across prompting-based, SFT (i.e., behavior cloning), and RL-based methods. We hypothesize that this form of test-time scaling facilitates deeper exploration and yields longer trajectories, potentially enabling the agent to iteratively refine its actions and make more informed decisions through extended interactions.
WebAgent-R1 generalizes well to out-of-distribution (OOD) tasks.
We conducted additional evaluation on the WebVoyager benchmark using Qwen2.5-3B as the baseline model. The benchmark covers diverse domains that are unseen in the WebArena environment and thus serves as an OOD evaluation of our method. Specifically, we randomly sample 25 tasks for each of 5 domains in WebVoyager as the OOD evaluation set to compare our method with both the prompting baseline and the SFT variant without any further training. As shown in Table 4, WebAgent-R1 consistently outperforms both prompting and SFT baselines across all domains, confirming the effectiveness and generalizability of our method.
Table 4: Out-of-distribution (OOD) evaluation on five domains from the WebVoyager benchmark. We report the success rates of compared methods on each domain.
Domain Prompting SFT WebAgent-R1
Allrecipes 0% 4% 28%
Amazon 4% 4% 24%
Arxiv 20% 20% 24%
Coursera 16% 16% 44%
Google Map 4% 16% 40%
Average 8.8% 12% 32%
@@ 4.1 LLM-based Agents
LLMs have demonstrated promising agentic capabilities, such as breaking down complex tasks into manageable subgoals and reasoning over long horizons (Zhou et al., 2022; Huang et al., 2022; Madaan et al., 2022; Li et al., 2023a, b; Wu et al., 2024; Liu et al., 2025b; Chu et al., 2025). Building on these capabilities, LLM-based agents have been applied to a variety of real-world interactive tasks, including web navigation (Nakano et al., 2021; Yao et al., 2022; Ma et al., 2023; Gur et al., 2024; Abuelsaad et al., 2024; Lutz et al., 2024; Patel et al., 2024; Putta et al., 2024), general computer use (Li et al., 2020; Deng et al., 2023; Yang et al., 2024), and embodied environments (Puig et al., 2018; Shridhar et al., 2020; Toyama et al., 2021; Fan et al., 2022; Huang et al., 2022). Specifically, our work focuses on text-based web agents that operate in browser-based environments purely based on HTML content, which requires agentic capabilities such as tool use, memory, and decision-making under partial observability (Zhou et al., 2024a; Qi et al., 2025). Complementary to this line of work, GUI agents leverage additional multimodal inputs such as screenshots, enabling visual-guided interactions with the environment (Lee et al., 2023; Shaw et al., 2023; Zheng et al., 2024; He et al., 2024a, b; Koh et al., 2024; Kil et al., 2024; Lei et al., 2025; Liu et al., 2025a). For a comprehensive overview, we refer readers to recent surveys (Wang et al., 2024a; Tseng et al., 2024; Hu et al., 2025; Ning et al., 2025).
@@ 4.2 Reinforcement Learning for LLMs
Recent advances like DeepSeek-R1 (Guo et al., 2025) highlight the strong potential of RL in enhancing LLMs. However, most prior work focuses on single-turn tasks such as math problems (Shao et al., 2024; Zhu et al., 2025; Shao et al., 2025; Ouyang et al., 2025; Wei et al., 2025), with limited exploration in multi-turn settings (Zhou et al., 2024b, 2025). Recent efforts have made some progress in this direction, such as training LLM agents to repeatedly use search engines (Jin et al., 2025; Sun et al., 2025; Chen et al., 2025; Song et al., 2025; Wang et al., 2025a), but typically constrain actions to simple API calls without real environment interaction. A few concurrent works, such as RAGEN (Wang et al., 2025b) and SkyRL (Cao et al., 2025), have applied RL to more dynamic settings like simulated games and coding environments (Jimenez et al., 2024). However, real-world web environments remain largely underexplored. Our work fills this gap by providing a practical framework and offering actionable insights for training web agents with end-to-end RL.
@@ 5 Conclusion
This work introduces WebAgent-R1, an end-to-end multi-turn RL framework for training web agents. We extend the standard GRPO to multi-turn settings, termed M-GRPO, and implement dynamic context compression and parallel trajectory rollout mechanisms for efficient training. Empirically, WebAgent-R1 achieves new state-of-the-art results on the WebArena-Lite benchmark. Our findings underscore the critical role of behavior cloning in initializing web agents, providing a strong foundation for effective RL. We further analyze training dynamics and explore the effects of thinking-based prompting and test-time scaling strategies, showing that increasing interaction depth consistently enhances web agents. Future work includes exploring multi-modal inputs and extending our approach to broader GUI-based tasks beyond web environments, such as computer use.
Limitations and Potential Risks
Despite the effectiveness of WebAgent-R1, our current approach has several limitations that suggest directions for future work. First, we consider only textual input for the web tasks. Incorporating additional visual input (e.g., screenshots) may enhance performance since visual information, such as layout and colors, can be helpful for effective navigation and decision-making. Second, our method relies on rule-based outcome rewards to guide RL training. While effective in our setting, such reward functions may not be readily available in other interactive scenarios, such as open-ended travel planner agents, where task goals are ambiguous and no clear reference or verifiable outcome is available. Lastly, like existing web agents, our model is trained with a fixed set of predefined actions (e.g., click, type), which can limit its flexibility when encountering interactive elements that require unseen operations. Enabling dynamic adaptation to new operations remains an open challenge for web agents.
In terms of potential risks, such agents should be used with caution when deployed in real-world environments, especially those involving administrative privileges. For example, when interacting with content management systems (CMS) in a production environment, the agent may inadvertently perform destructive actions, such as modifying or deleting sensitive data. To ensure safe deployment, future work should incorporate permission controls, verification prompts, and safeguards to prevent high-impact or irreversible actions.
Acknowledgments
The authors would like to thank Yu Meng and Shiyu Feng from the University of Virginia for their valuable feedback and discussions. We also thank anonymous reviewers for their constructive and insightful comments.